# A Copula Based Supervised Filter for Feature Selection in Diabetes Risk Prediction Using Machine Learning

### v2

@@ Abstract
Effective feature selection is critical for building robust and interpretable predictive models, particularly in medical applications where identifying risk factors in the most extreme patient strata is essential. Traditional methods often focus on average associations, potentially overlooking predictors whose importance is concentrated in the tails of the data distribution. In this study, we introduce a novel, computationally efficient supervised filter method that leverages the Gumbel copula’s upper-tail dependence coefficient ($\lambda_{U}$) to rank features based on their tendency to be simultaneously extreme with a positive outcome. We conducted a rigorous evaluation of this method against four standard baselines (Mutual Information, mRMR, ReliefF, and L1/Elastic-Net) using four distinct classifiers on two diabetes datasets: a large-scale public health survey (CDC, N=253,680) and a classic clinical benchmark (PIMA, N=768). Our analysis included comprehensive statistical tests, permutation importance, and robustness checks. On the CDC dataset, our method was the fastest selector and reduced the feature space by $\approx$ 52% while maintaining predictive performance statistically indistinguishable from a model using all features. On the PIMA dataset, our method’s feature ranking yielded the single best-performing model, achieving the highest ROC-AUC of all tested configurations. Across both datasets, the Gumbel-$\lambda_{U}$ selector consistently identified clinically coherent and impactful predictors. We conclude that feature selection via upper-tail dependence is a powerful, efficient, and interpretable new tool for developing risk models in public health and clinical medicine.
@@ Introduction
Machine learning has emerged as a powerful paradigm for developing such tools, leveraging large-scale health datasets like the CDC Diabetes Health Indicators survey to uncover complex patterns associated with disease risk [2]. A critical step in building effective and trustworthy predictive models is feature selection, a process fundamental to both performance and scientific utility [3]. For a non-specialist audience, feature selection (FS) is the principled process of identifying the most informative subset of predictors from a larger pool. The goals are threefold: (i) improve predictive performance by eliminating irrelevant or noisy variables that promote overfitting; (ii) reduce cost and latency by limiting variables to measure, store, and process; and (iii) enhance interpretability to yield actionable domain insight.
@@
We validate on two public datasets. On the large CDC survey, using just the top-10 Gumbel-ranked features (a 52% reduction from 21 to 10) and a Gradient Boosting classifier, we achieve ROC–AUC 0.823, statistically indistinguishable from using all 21 features (DeLong $p>0.99$ for all pairwise comparisons) while remaining highly robust to light perturbations and computationally fastest among tested selectors. On the clinical PIMA benchmark (8 predictors), the Gumbel ranking paired with Random Forest attains the top ROC–AUC 0.867, slightly surpassing strong baselines, and aligns with domain expectations (glucose, BMI, age dominate). Overall, our results show that focusing on the extremes can create models that are simple, easy to understand, and effective for clinical and public health use.
@@
Feature selection methods are usually grouped as filters, wrappers, or embedded approaches [3]. Wrappers use a learning algorithm to evaluate feature subsets, such as genetic search or RFE, and often achieve strong results but require significant computation [12, 13, 14]. Embedded methods combine selection with model training. Well-known examples include LASSO and elastic-net, which encourage sparsity through regularization [15, 16, 17].
Our work introduces a filter method where features are scored directly from the data, without relying on any downstream model [18]. Classic filters include mutual information (MI), which measures non-linear dependence but evaluates features one at a time [5]. The mRMR method addresses this by selecting features that are both highly relevant and minimally redundant [19, 20]. ReliefF focuses on local class separation by using nearest neighbors [21, 22]. However, these methods still mainly assess overall association. Unlike correlation-based filters (which are tail-independent), our score targets joint extremes via the Gumbel upper-tail coefficient $\lambda_{U}$ rather than average association.
A recent area of research uses copulas to create more fair and robust ways to measure dependence. Copula-based filters like Copula Correlation (Ccor) and Robust Copula Dependence (RCD) help reduce functional bias compared to traditional metrics [23, 24], which aligns with the goals of the MIC literature [25, 26]. Other advances include using copulas for dimensionality reduction [27], entropy-based selection [28, 29], hybrid filters that combine copula and information-theoretic scores. However, most copula-based studies still focus on measuring overall dependence or redundancy between features.
Importantly, tail-dependence coefficients, which directly measure the probability of joint extremes, have not yet been used as the main supervised ranking criterion in feature selection. Peng et al. have pointed out that a Gaussian copula cannot capture tail dependence and have called for methods that use tail-dependent families. Our work addresses this need. By using the Gumbel copula’s upper-tail coefficient $\lambda_{U}$ [11, 7], we shift the selection focus from overall association to joint extremal behavior. This approach provides a simple, interpretable, and computationally efficient supervised filter that (i) identifies clinically meaningful predictors linked to high-risk cases, (ii) matches top performance on both large survey and clinical datasets, and (iii) remains fast compared to standard baselines like MI, mRMR, and ReliefF, making it practical for real-world use.
@@
A copula $C:[0,1]^{2}\to[0,1]$ [10, 11] is a joint distribution with uniform marginals that captures dependence independently of the marginal distributions. By Sklar’s theorem [9], for any bivariate cdf $H$ with continuous marginals $F_{X},F_{Y}$ there exists a unique copula $C$ such that
@@
How we use $\lambda_{U}$ for feature selection.
@@
Our proposed Gumbel-$\lambda_{U}$ procedure is a supervised filter method. In comparisons, MI, mRMR, and ReliefF are also filters; L1/Elastic-Net is an embedded selector.
@@
MI [4, 5] is a filter method that measures how far the joint behavior of two variables deviates from what we would see if they were independent. If $X$ and $Y$ were independent, the joint distribution would factor as $p(x,y)=p(x)p(y)$. MI compares the real $p(x,y)$ to that independence baseline: bigger differences $\Rightarrow$ stronger dependence (linear or non-linear).
@@
$I(X;Y)\geq 0$, and $I(X;Y)=0$ iff $X$ and $Y$ are independent.
Use in Feature Selection: Rank features by $I(X_{j};Y)$.
@@
mRMR [19, 20] is a filter method that builds a set step-by-step. At each step, it prefers a feature that is (i) relevant to the label (informative on its own) and (ii) not redundant with features already chosen (brings new information). A common score is “MI with YY” minus “average similarity to selected features.” In practice, similarity is often measured by absolute correlation for speed.
Use in Feature Selection: Build the set greedily by adding the feature with the highest “relevance – redundancy” score, i.e., maximize $I(X_{j};Y)$ minus its average correlation with already-selected features; stop after $k$.
@@
ReliefF [21, 22] is a nearest-neighbor view of usefulness. It is a filter method. For each data point, it looks at the closest point in the same class (a “hit”) and the closest point in the other class (a “miss”). A feature gets a higher score if, along that feature, the point is close to its hit but far from its miss, i.e., the feature separates classes in local neighborhoods. This captures non-linear structure and some interactions.
Use in Feature Selection: For many sampled points, compare each feature’s average difference to nearest neighbors of the other class vs the same class; rank/select by the resulting average score (bigger = more discriminative).
@@
L1EN [16, 15, 17] is an embedded method.
What are the penalties?
L1 (lasso) penalty: sum of absolute values $\|\beta\|_{1}=\sum_{j}|\beta_{j}|$. Tends to set many coefficients exactly to zero $\Rightarrow$ sparse model $\Rightarrow$ selected features are those with non-zero coefficients.
L2 (ridge) penalty: sum of squares $\|\beta\|_{2}^{2}=\sum_{j}\beta_{j}^{2}$. Shrinks coefficients toward zero but rarely makes them exactly zero $\Rightarrow$ stabilizes estimates when features are correlated.
Elastic-Net combines both: $\alpha\|\beta\|_{1}+\frac{1-\alpha}{2}\|\beta\|_{2}^{2}$, giving sparsity (from L1) and stability (from L2).
We fit a logistic model and let the penalty choose which coefficients should be non-zero. Non-zero $\beta_{j}$ means feature $j$ is “kept”; zero means it’s effectively dropped.
Use in Feature Selection: Fit a penalized logistic regression on the training set and rank/select features by the absolute value of their learned coefficients (larger $|\beta_{j}|$ $\Rightarrow$ more important; pick top-$k$).
@@
RF [34] creates many decision trees using different random samples of the data and by choosing random subsets of features for each split. The final prediction is made by combining the results from all the trees, either by voting or averaging.
It works for both classification and regression. Low variance from averaging, handles non-linearities and interactions, and works well on tabular data with mixed signals.
@@
GB [35] builds a series of trees, where each new tree tries to correct the mistakes of the previous ones. The final model adds together the results of many small trees.
It works for both classification and regression. Strong accuracy on tabular data, with fine control of overfitting via learning rate and tree depth.
@@
XGB [36] is a high-performance, regularized implementation of gradient boosting with system optimizations (parallelism, sparsity-aware learning).
It works for both classification and regression. Typically fast and accurate on medium/large tabular problems; regularization helps when features are correlated.
@@
LR [37], despite its name, is mainly used for classification tasks, either with two classes or more. It calculates a score from the features and then uses a logistic or softmax function to turn this score into class probabilities.
It is simple, interpretable, often well-calibrated probabilities; with regularization it resists overfitting and can act as an embedded selector.
@@
Thus, ROC–AUC jointly refers to the ROC curve and its area. Values near 1 indicate strong ranking ability; 0.5 corresponds to random guessing. Because it does not depend on a single threshold (and is less sensitive to class imbalance than accuracy), ROC–AUC is our primary metric.
@@
A high p-value (e.g., p ¿ 0.05) suggests that any observed difference in AUC is not statistically significant.
@@
Dataset Description (CDC)
@@
Demographic Variables Age, Gender, Income, Education Level
@@
Target Variable: Outcome is binary, where 0 denotes no diabetes and 1 denotes diabetes. Our study treats risk prediction as a binary classification task, consistent with the dataset’s label.
@@
This section describes the entire pipeline we use in the paper. It covers data preparation, feature selection, model training, thresholding, and statistical testing. The implementation is deterministic with a fixed seed of 123. Additionally, all splits are stratified to maintain class balance.
@@
For each candidate predictor $X_{j}$, we ask whether the upper-tail dependence coefficient $\lambda_{U}(X_{j},Y)$ is materially greater than zero and whether ranking features by $\lambda_{U}$ yields classifiers with higher ROC-AUC (and related metrics) than baselines (MI, mRMR, ReliefF, and L1EN). Formally, for screening we consider
$H_{0}:\lambda_{U}(X_{j},Y)=0\text{ vs. }H_{1}:\lambda_{U}(X_{j},Y)>0,$ (15)
and for prediction we compare models trained on the $\lambda_{U}$-selected set against baselines using paired tests (DeLong for AUC; McNemar for error).
Guided by the research question above, we compute for each feature $X_{j}$ the Gumbel-based upper-tail dependence score $\lambda_{U}(X_{j},Y)$ from pseudo-observations. We rank features by $\lambda_{U}$, fix $k$ per dataset (CDC: k=10k{=}10; PIMA: k=8k{=}8), and evaluate the resulting sets alongside baselines (MI, mRMR, ReliefF, L1EN) using four classifiers (RF, XGB, LR, GB) with class-balanced training and probability calibration. We report Accuracy, Precision, Recall, F1, and ROC-AUC; we compute permutation importance on the held-out test set; and we use DeLong and McNemar tests for paired comparison. Computationally, computing Kendall’s $\tau$ per feature with a closed-form transform yields an $O(d\,n\log n)$ selector with no model fitting in the selection loop.
We study two public datasets: CDC dataset and the PIMA Indians Diabetes dataset. For CDC, we load the features and convert the target into binary. For PIMA, we use the common practice of treating zeros as missing for five physiological variables: Glucose, BloodPressure, SkinThickness, Insulin, and BMI. In the training split only, we replace zeros in these columns with NaN and fit a median imputer. We then apply the learned medians to both the training and test splits. No other imputations or transformations are applied at the dataset level.
We split each dataset once into training and test sets using an 80/20 stratified split. From the training portion, we hold out a validation subset, which is 20% of the training data. This subset is used only to select decision thresholds; we never tune models on the test set. We address class imbalance in two ways. First, when training logistic regression, random forests, and gradient boosting, we set class weights to “balanced.” For XGBoost, we calculate scale_pos_weight as the ratio of negatives to positives in the training data. Second, instead of fixing the probability cutoff at 0.5, we choose the operating threshold on the validation data by testing 199 quantiles of the predicted probabilities. We then select the threshold that maximizes the F1 score. This approach keeps operating points comparable across different class prevalences and reduces the risk that an imbalanced base rate affects the differences in accuracy.
@@
    The elastic-net baseline fits a logistic regression with l1_ratio = 0.7 using the SAGA solver on standardized predictors. It runs for up to 4,000 iterations with
    class_weight=‘‘balanced’’ and ranks features by the absolute coefficient magnitudes at the end.
@@
    The mRMR implementation is a simple greedy method that selects the first feature with the highest mutual information to the label (using the same estimator as before). It then adds the feature that maximizes “relevance minus redundancy,” where redundancy is the mean absolute Pearson correlation with the already chosen features.
@@
Our method is a supervised filter feature selector. It is supervised because the score for each feature $X_{j}$ uses the label $Y$ (via $\lambda_{U}(X_{j},Y)$). It is a filter because features are scored and ranked without fitting any predictive model during selection; the score depends only on the data’s dependence structure (via pseudo-observations and the Gumbel mapping), not on a particular classifier. By contrast, wrappers (e.g., forward selection, RFE) repeatedly train a model to evaluate subsets, and embedded methods (e.g., L1/Elastic-Net) perform selection inside model training.
@@
For the CDC dataset, we use the top-10 Gumbel-$\lambda_{U}$ features. For PIMA, we use the top-8, which matches the full set because PIMA has eight covariates. We picked $k$ by briefly exploring candidate values (CDC: $k\in\{3,5,7,10\}$; PIMA: $k\in\{3,5,8\}$). For each $k$, we ranked features by Gumbel-$\lambda_{U}$ on the training split. We trained the same models, adjusted the decision threshold on a validation split using the F1 score, and compared performance using ROC-AUC as the main metric given its threshold-free nature and robustness under class imbalance. This process led us to select $k=10$ for CDC and $k=8$ for PIMA. We then kept these values constant for all analyses in the paper.
We evaluate a diverse panel of classifiers on each feature set:
  1. 1.
    Logistic regression (LR): Maximum 2,000 iterations, class_weight=‘‘balanced’’.
  2. 2.
    Random forests (RF): 400 trees, class_weight=‘‘balanced’’.
  3. 3.
    Gradient boosting (GB): Scikit-learn’s gradient boosting with a fixed random seed.
  4. 4.
    XGBoost (XGB): 400 trees, maximum depth of 6, learning rate of 0.05, subsample of 0.9, colsample_bytree of 0.9, histogram tree method, and scale_pos_weight as described below.
To address class imbalance during model training, we incorporate specific weighting strategies. For LR and RF, the class_weight=’balanced’ parameter is used, which automatically adjusts weights to be inversely proportional to the class frequencies present in the training data. For GB, we emulate class balancing by passing per-sample weights proportional to the inverse class frequencies. For XGB, we set the scale_pos_weight parameter equal to the ratio of the number of negative samples to positive samples.
@@
    The third is MCAR missingness (missing completely at random), where we randomly set 10% of entries to missing per feature and then fill in those gaps using the training median for that feature.
We evaluate each change using the same training process, and we record ROC–AUC scores on the unchanged test labels.
@@
9:   Build model. Use class balancing (class_weight=’balanced’ for RF/LR/GB; scale_pos_weight for XGB). Apply isotonic calibration (5-fold) for RF/XGB/GB. Standardize inputs for LR only.
@@
11:   Final fit & test. Refit on full train with $S$, apply the chosen threshold to test scores; record ROC–AUC, accuracy, precision, recall, F1; save ROC/PR curves and confusion matrix.
@@
14:Pick best configuration. By primary metric (ROC–AUC), identify the best (feature set, model).
@@
  • Compute PI on the test set for the overall best (set, model) using permutation_importance with scoring=’roc_auc’, n_repeats=500n\_repeats{=}500.
@@
17:Significance tests. For the best model, compare the Gumbel set against other sets using paired DeLong tests on ROC–AUC and McNemar’s test on correctness.

### v4

# A Copula Based Supervised Filter for Feature Selection in Machine Learning Driven Diabetes Risk Prediction

@@ Abstract
Effective feature selection is critical for building robust and interpretable predictive models, particularly in medical applications where identifying risk factors in the most extreme patient strata is essential. Traditional methods often focus on average associations, potentially overlooking predictors whose importance is concentrated in the tails of the data distribution. In this study, we introduce a novel, computationally efficient supervised filter that leverages the Gumbel copula’s upper-tail dependence coefficient ($\lambda_{U}$) to rank features by their tendency to be simultaneously extreme with the positive class. We evaluated this method against four standard baselines (Mutual Information, mRMR, ReliefF, and L1/Elastic-Net) across four classifiers on two diabetes datasets: a large-scale public health survey (CDC, N=253,680N{=}253{,}680) and a classic clinical benchmark (PIMA, N=768N{=}768). Our analysis included comprehensive statistical tests, permutation importance, and robustness checks. On the CDC dataset, our method was the fastest selector and reduced the feature space by $\approx$52%. While this resulted in a minimal but statistically significant performance trade-off compared to using all 21 features, our filter significantly outperformed standard filters (Mutual Information, mRMR) and was statistically indistinguishable from the strong ReliefF baseline. On the PIMA dataset (8 predictors), our method’s ranking produced the numerically highest ROC-AUC, despite paired DeLong tests showing no statistically significant differences versus strong baselines. PIMA thus serves as a ranking-only sanity check that our upper-tail criterion behaves sensibly in a low-dimensional clinical setting. Across both datasets, the Gumbel-$\lambda_{U}$ selector consistently identified clinically coherent and impactful predictors. We conclude that feature selection via upper-tail dependence is a powerful, efficient, and interpretable tool for developing risk models in public health and clinical medicine.
@@ Introduction
Machine learning has emerged as a powerful paradigm for developing such tools, leveraging large-scale health datasets like the CDC Diabetes Health Indicators[2] survey to uncover complex patterns associated with disease risk. A critical step in building effective and trustworthy predictive models is feature selection, a process fundamental to both performance and scientific utility [3]. For a non-specialist audience, feature selection (FS) is the principled process of identifying the most informative subset of predictors from a larger pool. The goals are threefold: (i) improve predictive performance by eliminating irrelevant or noisy variables that promote overfitting; (ii) reduce cost and latency by limiting variables to measure, store, and process; and (iii) enhance interpretability to yield actionable domain insight.
@@
We validate on two public datasets. On the large CDC survey, our method demonstrates a powerful trade-off between parsimony and performance. By reducing the feature set by 52% (from 21 to 10), our Gumbel-selected model achieves a competitive ROC-AUC of 0.823 and is the fastest selector. Most importantly, it proves its efficacy as a top-tier filter by statistically outperforming standard filter methods like Mutual Information and mRMR, while performing on par with the strong ReliefF baseline. To demonstrate external validity beyond a large, survey-style cohort (CDC), we also evaluate on the PIMA Indians Diabetes dataset, a small clinical benchmark with eight predictors. Because PIMA contains all and only eight covariates, this experiment isolates feature ranking rather than dimensionality reduction: all methods use the same variables. The Gumbel ranking combined with a Random Forest attains the highest ROC-AUC of 0.867, with paired DeLong tests showing no significant differences across selectors (DeLong $p>0.05$). As such, PIMA functions as a ranking-only sanity check that our upper-tail criterion yields clinically coherent orderings without harming discrimination. Overall, our results show that focusing on the extremes can create models that are simple, easy to understand, and effective for clinical and public health use.
@@
Feature selection methods are usually grouped as filters, wrappers, or embedded approaches [3]. Wrappers[12] use a learning algorithm to evaluate feature subsets, such as genetic search or RFE, and often achieve strong results but require significant computation [13, 14]. Embedded methods combine selection with model training. Well-known examples include LASSO[15] and elastic-net, which encourage sparsity through regularization [16, 17].
Our work introduces a filter method where features are scored directly from the data, without relying on any downstream model [18]. Classic filters include mutual information (MI), which measures non-linear dependence but evaluates features one at a time [5]. The mRMR method addresses this by selecting features that are both highly relevant and minimally redundant [19, 20]. ReliefF[21, 22] focuses on local class separation by using nearest neighbors . However, these methods still mainly assess overall association. Unlike correlation-based filters (which are tail-independent), our score targets joint extremes via the Gumbel upper-tail coefficient $\lambda_{U}$ rather than average association.
A recent area of research uses copulas to create more fair and robust ways to measure dependence. Copula-based filters like Copula Correlation (Ccor) and Robust Copula Dependence (RCD) help reduce functional bias compared to traditional metrics [23, 24], which aligns with the goals of the MIC literature [25, 26]. Other advances include using copulas for dimensionality reduction [27], entropy-based selection [28, 29], hybrid filters that combine copula and information-theoretic scores to balance relevance and redundancy[30]. However, most copula-based studies still focus on measuring overall dependence or redundancy between features.
Importantly, tail-dependence coefficients, which directly measure the probability of joint extremes, have not yet been used as the main supervised ranking criterion in feature selection. As seen from the literature, a Gaussian copula cannot capture tail dependence and have called for methods that use tail-dependent families. Our work addresses this need. By using the Gumbel copula’s upper-tail coefficient $\lambda_{U}$ [11, 7], we shift the selection focus from overall association to joint extremal behavior. This approach provides a simple, interpretable, and computationally efficient supervised filter that (i) identifies clinically meaningful predictors linked to high-risk cases, (ii) matches top performance on both large survey and clinical datasets, and (iii) remains fast compared to standard baselines like MI, mRMR, and ReliefF, making it practical for real-world use.
Finally, to specifically assess the effect of tail-aware ranking when dimensionality reduction is moot, we include the 8-variable PIMA benchmark where all methods use the same variables.
@@
A copula $C:[0,1]^{2}\to[0,1]$ is a joint distribution with uniform marginals that captures dependence independently of the marginal distributions[10, 11]. By Sklar’s theorem [9], for any bivariate cdf $H$ with continuous marginals $F_{X},F_{Y}$ there exists a unique copula $C$ such that
@@
Feature Ranking via $\lambda_{U}$.
@@
Our proposed Gumbel-$\lambda_{U}$ procedure is a supervised filter method. In comparisons, MI, mRMR, and ReliefF are also filters; L1/Elastic-Net is an embedded selector. Benchmarking the proposed method against wrapper and hybrid methods will be left for future work.
@@
MI [4, 5] is a filter method that measures how far the joint behavior of two variables deviates from what would be expected if they were independent. If $X$ and $Y$ were independent, the joint distribution would factor as $p(x,y)=p(x)p(y)$. MI quantifies how the actual joint relationship between two variables differs from this independence baseline; a greater deviation indicates a stronger dependence, whether linear or non-linear.
@@
$I(X;Y)\geq 0$, and $I(X;Y)=0$ if and only if $X$ and $Y$ are independent. In feature selection, MI is used to rank features by their dependence with the target variable, typically by computing $I(X_{j};Y)$ for each feature $X_{j}$.
@@
mRMR [19, 20] is a filter method that builds a feature set step-by-step. At each step, it prefers a feature that is (i) relevant to the label (informative on its own) and (ii) not redundant with features already chosen (brings new information). A common score is the difference between “MI with YY” and the “average similarity to selected features”. In practice, similarity is often measured by absolute correlation for computational efficiency. In feature selection, mRMR builds the feature set greedily by adding the feature with the highest “relevance – redundancy” score, i.e., maximizing $I(X_{j};Y)$ minus its average correlation with already-selected features, and stops after selecting $k$ features.
@@
ReliefF [21, 22] is a filter method that takes a nearest-neighbor view of feature usefulness. For each data point, it finds the closest point in the same class (a “hit”) and the closest point in the other class (a “miss”). A feature receives a higher score if, along that feature, the point is close to its hit but far from its miss, meaning the feature helps separate classes in local neighborhoods. This approach captures non-linear structure and some feature interactions. In feature selection, ReliefF evaluates many sampled points by comparing each feature’s average difference to nearest neighbors of the other class with that of the same class. Features are then ranked or selected based on the resulting average score. Larger scores indicate more discriminative features.
@@
L1EN [16, 15, 17] is an embedded method that uses both $L1$ and $L2$ regularization penalties, with the emphasis on $L1$ for feature selection. The $L1$ (lasso) penalty, $\|\beta\|_{1}=\sum_{j}|\beta_{j}|$, tends to set many coefficients exactly to zero, producing a sparse model where selected features correspond to the non-zero coefficients. The $L2$ (ridge) penalty, $\|\beta\|_{2}^{2}=\sum_{j}\beta_{j}^{2}$, shrinks coefficients toward zero but rarely makes them exactly zero, stabilizing estimates when features are correlated. Elastic Net combines both penalties, $\alpha\|\beta\|_{1}+\frac{1-\alpha}{2}\|\beta\|_{2}^{2}$, providing sparsity from $L1$ and stability from $L2$. When a penalized logistic regression is fitted with these penalties, the $L1$ component determines which coefficients are exactly zero, with a non-zero coefficient $\beta_{j}$ indicating that feature $j$ is retained and a zero coefficient effectively removing it. Feature selection can then be performed by ranking features according to the absolute values of their coefficients, $|\beta_{j}|$, with larger values indicating greater importance, and selecting the top $k$ features based on this ranking.
@@
RF [35] creates many decision trees using different random samples of the data and by choosing random subsets of features for each split. The final prediction is made by combining the results from all the trees, either by voting or averaging. RF is effective for both classification and regression tasks. It reduces variance through averaging, handles non-linearities and feature interactions, and performs well on tabular data with mixed signals.
@@
GB [36] constructs a sequence of trees, with each new tree trained to correct the errors of the preceding ones. The final model aggregates the predictions of all trees. GB is applicable to both classification and regression tasks. It offers strong predictive performance on tabular data, with fine control over overfitting through the learning rate and tree depth.
@@
XGB [37] is a high-performance, regularized implementation of gradient boosting that incorporates system-level optimizations such as parallelism and sparsity-aware learning. It is typically fast and accurate on medium to large tabular datasets, and its regularization mechanisms help mitigate issues when features are correlated.
@@
LR [38], despite its name, is primarily used for classification tasks with two or more classes. It computes a linear combination of the features and applies a logistic or softmax function to convert this score into class probabilities. LR is simple and interpretable, often provides well-calibrated probabilities, and with regularization, it resists overfitting and can serve as an embedded feature selector.
@@
Thus, ROC-AUC jointly refers to the ROC curve and its area. Values near 1 indicate strong ranking ability; 0.5 corresponds to random guessing. Because it does not depend on a single threshold (and is less sensitive to class imbalance than accuracy), ROC-AUC is our primary metric.
@@
A high p-value (e.g., $p>0.05$) suggests that any observed difference in AUC is not statistically significant. It does not give the direction. Direction comes from the numerical AUCs themselves.
@@
Dataset Description (CDC Diabetes Health Indicators)
@@
Demographic Variables Age, Sex, Income, Education Level
@@
Target Variable: The dataset includes a binary label Outcome (0 = no diabetes, 1 = diabetes). We use this label as provided by the repository—no relabeling or class collapsing was performed. Our study treats diabetes risk prediction as a binary classification problem, in line with the repository label. This definition enables focused evaluation of risk factors and classification performance.
@@
Note on scope. Since PIMA has exactly eight predictors, all methods ultimately use the same variables. We report rankings only for interpretability; statistical tests (paired DeLong and McNemar) verify that any observed performance differences are not significant.
@@
For tree ensembles (RF/GB/XGB), predicted probabilities were calibrated via 5-fold isotonic regression [42, 43, 44]. Logistic regression was left uncalibrated. Calibration was fit on the training folds only and applied to the validation/test sets. We calibrated the tree-ensemble probabilities because uncalibrated trees are typically miscalibrated (often overconfident), whereas regularized logistic regression is generally close to calibrated; therefore, we left LR uncalibrated. Logistic regression used standardized features, while tree-based models used raw inputs. Decision thresholds were selected on the validation set to maximize F1 and then held fixed for test evaluation. All experiments used a fixed random seed and identical train/validation/test splits.
We standardize inputs only where the downstream method requires distance or coefficient comparability:(i) LR is trained on standardized predictors (scaler fit on train only and applied to val/test) to stabilize the optimizer and penalty. (ii) ReliefF relies on Euclidean nearest neighbors, so we z-score each feature on the training split only before its hit/miss computations and apply the learned mean/SD to val/test. By contrast, (iii) tree ensembles (RF/GB/XGB) are scale-invariant and use raw inputs; and (iv) the MI and mRMR filters are used without scaling: MI here is univariate $I(X_{j};Y)$, so linear rescaling of $X_{j}$ does not change the information content, and the mRMR redundancy term uses Pearson correlation, which is itself scale-invariant. All scalers are train-only to avoid leakage.
@@
For each feature $X_{j}$, we compute the Gumbel upper–tail dependence score $\lambda_{U}(X_{j},Y)$ from rank-based pseudo-observations and rank features by $\lambda_{U}$. We say a feature shows a positive upper-tail association when the 95% bootstrap confidence interval for $\lambda_{U}$ lies above 0. We keep the top-$k$ features and train models on that set. To judge usefulness, we compare these models with baselines (MI, mRMR, ReliefF, and L1/Elastic-Net) on the same test split, using ROC-AUC as the primary metric (with PR–AUC, accuracy, recall, and F1 as secondary metrics). For formal paired comparisons we use DeLong’s test for ROC-AUC and McNemar’s test for error profiles.
Guided by the research question above, we rank features by $\lambda_{U}$, fix $k$ per dataset (CDC: k=10k{=}10; PIMA: k=8k{=}8), and evaluate the resulting sets alongside baselines (MI, mRMR, ReliefF, L1EN) using four classifiers (RF, XGB, LR, GB) with class-balanced training and probability calibration. We report Accuracy, Precision, Recall, F1, and ROC-AUC; we compute permutation importance on the held-out test set; and we use DeLong and McNemar tests for paired comparison. Computationally, computing Kendall’s $\tau$ per feature with a closed-form transform yields an $O(d\,n\log n)$ selector with no model fitting in the selection loop.
We study two public datasets: CDC Diabetes Health Indicators dataset and the PIMA Indians Diabetes dataset. For CDC, we use all 21 features as provided. The dataset has no missing values. For PIMA, we use the common practice of treating zeros as missing for five physiological variables: Glucose, BloodPressure, SkinThickness, Insulin, and BMI. In the training split only, we replace zeros in these columns with NaN and fit a median imputer. We then apply the learned medians to both the training and test splits.
We split each dataset once into training and test sets using an 80/20 stratified split. From the training portion, we hold out a validation subset, which is 20% of the training data. This subset is used only to select decision thresholds; we never tune models on the test set. We address class imbalance in two ways. First, when training logistic regression, random forests, and gradient boosting, we set class weights to “balanced”. For XGBoost, we calculate scale_pos_weight as the ratio of negatives to positives in the training data. Second, instead of fixing the probability cutoff at 0.5, we choose the operating threshold on the validation data by scanning a fine grid of probability quantiles and selecting the cutoff that maximizes the F1 score. This approach keeps operating points comparable across different class prevalences and reduces the risk that an imbalanced base rate affects the differences in accuracy.
@@
    The elastic-net baseline fits a logistic regression using the SAGA solver on standardized predictors. It runs for up to 4,000 iterations with class_weight=‘‘balanced’’ and ranks features by the absolute coefficient magnitudes at the end.
@@
    The mRMR implementation is a simple greedy method that selects the first feature with the highest mutual information to the label (using the same estimator as before). It then adds the feature that maximizes “relevance minus redundancy”, where redundancy is the mean absolute Pearson correlation with the already chosen features.
@@
Our method is a supervised filter feature selector. It is supervised because the score for each feature $X_{j}$ uses the label $Y$ (via $\lambda_{U}(X_{j},Y)$). It is a filter because features are scored and ranked without fitting any predictive model during selection; the score depends only on the data’s dependence structure (via pseudo-observations and the Gumbel mapping), not on a particular classifier. Because the score is computed from pseudo-observations (ranks), it is invariant to monotone rescalings of $X_{j}$ and does not assume common measurement units. By contrast, wrappers (e.g., forward selection, RFE) repeatedly train a model to evaluate subsets, and embedded methods (e.g., L1/Elastic-Net) perform selection inside model training.
@@
For the CDC dataset, we chose the number of features via a small sweep, $k\in\{3,5,7,10\}$, and fixed $k=10$ for all CDC analyses. For PIMA, no $k$-sweep was performed: the dataset has exactly eight predictors, so all methods use the same eight variables and our criterion serves as a ranking-only check. Models were trained on the training split, the operating threshold was selected on a validation split using F1, and test performance was summarized by ROC-AUC (primary) and related metrics. See Aich (2025) [45] for additional CDC sweep details.
We evaluate a diverse set of classifiers on each feature set. Logistic regression (LR) is used with balanced class weights, while random forests (RF) are also applied with balanced class weights. Gradient boosting (GB) is implemented using the standard scikit-learn version with a fixed random seed, and XGBoost (XGB) is employed with its standard setup, handling class weights as described below.
To address class imbalance during model training, we incorporate specific weighting strategies. For LR and RF, the class_weight=‘‘balanced’’ parameter is used, which automatically adjusts weights to be inversely proportional to the class frequencies present in the training data. For GB, we emulate class balancing by passing per-sample weights proportional to the inverse class frequencies. For XGB, we set the scale_pos_weight parameter equal to the ratio of the number of negative samples to positive samples.
@@
Ordering for “All” vs. ranked sets.
For CDC, “All” uses every feature. For PIMA, all methods use the same eight predictors; column order has no effect for LR/RF/GB/XGB. We include rankings only for interpretability; as expected, AUCs are statistically indistinguishable across selectors (DeLong p ¿ 0.05).
@@
    The third is MCAR missingness (missing completely at random), where we randomly set 10%, 20%, and 30% of the entries to missing per feature and imputed them using the training median of that feature. In this paper, we report results only for the 10% missingness scenario, while the 20% and 30% results are provided in the dissertation[45].
We evaluate each change using the same training process, and we record ROC-AUC scores on the unchanged test labels.
@@
9:   Build model. Use class balancing (class_weight=‘balanced’ for RF/LR/GB; scale_pos_weight for XGB). Apply isotonic calibration (5-fold) for RF/XGB/GB. Standardize inputs for LR only.
@@
11:   Final fit & test. Refit on full train with $S$, apply the chosen threshold to test scores; record ROC-AUC, accuracy, precision, recall, F1; save ROC/PR curves and confusion matrix.
@@
14:Pick best configuration. By primary metric (ROC-AUC), identify the best (feature set, model).
@@
  • Compute PI on the test set for the overall best (set, model) using permutation_importance with scoring=‘roc_auc’, n_repeats=500n\_repeats{=}500.
@@
17:Significance tests. For the best model, compare the Gumbel set against other sets using paired DeLong tests on ROC-AUC and McNemar’s test on correctness.